NvidiaIntely (nvidiaintely):

What is the average binding energy per nucleon for a C-12 nucleus with a mass defect of 0.0993 amu? (1 amu= 1.66 x 10-27 kg; 1 J = 1 kg m2/s2) 1.48 x 10-11 J 1.24 x 10-12 J 4.12 x 10-21 J 8.94 x 1015 J

OpenStudy (kevin):

E = m x c^2 =0.0993* (3*10^8)^2 = 8937000000000000 (D)

NvidiaIntely (nvidiaintely):

Okay, thats what I got again haha.

NvidiaIntely (nvidiaintely):

Thanks Kevin!

OpenStudy (kevin):

XD

OpenStudy (kevin):

u are welcome

NvidiaIntely (nvidiaintely):

oops wait its, 1.24 x 10-12 J

OpenStudy (kevin):

0_O

NvidiaIntely (nvidiaintely):

Response Feedback: Using the formula, E = mc2, substitute in the value provided. First, convert amu to kg: 0.0993 amu x 1.66 x 10-27 kg /amu = 1.648 x 10-28 kg. Then, substitute into E = mc2: E = (1.648 x 10-28 kg) (3.00 x 108 m/s)2 = 1.484 x 10-11 kg m2/s2 = 1.484 x 10-11 J. In the C-12 nucleus, there are 12 nucleons. To calculate the energy per nucleon, divide the total binding energy by the number of nucleons: (1.484 x 10-11 J)/ 12 = 1.24 x 10-12 J

OpenStudy (kevin):

Owh yeah.. we should convert it too kg

OpenStudy (kevin):

Sorry....

NvidiaIntely (nvidiaintely):

Oops, ow well. Now we know, its fine.

OpenStudy (kevin):

(1 amu= 1.66 x 10-27 kg)